OpenStudy (riley):

Calculate the dy/dx. You need not expand your answers. Help with more derivatives please!! >_< y = (8.43x^-0.1 - 0.5x^-1) / (3.4+x^2.2)

OpenStudy (riley):

I know to use the derivative quotient rule, but every time I work through this problem, I keep coming up with the wrong answer. So I don't know if I'm plugging it in wrong, or if I'm messing up on the simple math part.

OpenStudy (riley):

The decimal powers are laughing at me . . .

OpenStudy (riley):

with their shiny negative symbols . .

OpenStudy (riley):

yesh, that is how I started mine out O_O

OpenStudy (riley):

Oh man, mine didn't come out like that . . . wait I'm confusing myself.

OpenStudy (riley):

oh wait nvm i'm so dumb, you hadn't solve it through yet. But yes thats what I did!! I hope I did the math right.

OpenStudy (riley):

(-.843x+.5x^-2)(3.4+x^2.2) - (8.43x^-0.1 - 0.5x^-1)(2.2x^1.2) / (3.4+x^2.2)^2

OpenStudy (riley):

sorry mine doesn't look as neat as yours, thats how it came out when I solved through the math

OpenStudy (riley):

oof -.843x^-1.1*****

OpenStudy (riley):

all these numbers confuse me so badly, it's hard to keep up with it all >_<

OpenStudy (riley):

k now I'm stuck Q_Q

OpenStudy (riley):

Help . . .

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{(8.43(-0.1)x^{-0.1-1}-0.5(-1)x^{-1-1})(3.4-x^{2.2})-(0+2.2x^{2.2-1})(8.43x^{-0.1}-0.5x^{-1})}{(3.4+x^{2.2})^2} \] \[\frac{dy}{dx}=\frac{(-0.843x^{-1.1}+0.5x^{-2})(3.4+x^{2.2})-(0+2.2x^{1.2})(8.43x^{-0.1}-0.5x^{-1})}{(3.4+x^{2.2})^2}\]

OpenStudy (riley):

Yeah, I've gotten that far; but I'm not sure what to do after that.

myininaya (myininaya):

what do you want to do?

OpenStudy (riley):

Solve it so I don't have to look at the problem anymore ever again!!

myininaya (myininaya):

well we already found y'

myininaya (myininaya):

does your teacher want you to distribute on top or something and the combine any like terms if any exist?

OpenStudy (riley):

But don't I have to simplify?? O_O

myininaya (myininaya):

if it says to simplify then you have to find a way to do that but if it doesn't say simplify you are done

OpenStudy (riley):

I know I don't have to simplify the denomiinator, but I have to combine like terms I guess for the numerator.

OpenStudy (riley):

denominator*** sorry

OpenStudy (riley):

nevermind, its the stupid like terms and math where I keep messing up, the supposedly easy part. I will just try to figure it out again later I guess.

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{(-0.843x^{-1.1}+0.5x^{-2})(3.4+x^{2.2})-(0+2.2x^{1.2})(8.43x^{-0.1}-0.5x^{-1})}{(3.4+x^{2.2})^2}\] ------------------------------ \[(-0.843x^{-1.1}+0.5x^{-2})(3.4+x^{2.2})\] \[(-0.843x^{-1.1}+0.5x^{-2})(3.4)+(-0.843x^{-1.1}+0.5x^{-2})(x^{2.2})\] \[-0.28662x^{-1.1}+1.7x^{-2}+(-0.843x^{-1.1+2.2})+0.5x^{-2+2.2}\]

myininaya (myininaya):

----------------------------- \[(0+2.2x^{1.2})(8.43x^{-0.1}-0.5x^{-1})\] \[(2.2x^{1.2})(8.43x^{-0.1}-0.5x^{-1})\] \[18.546x^{1.2-0.1}-1.1x^{1.2-1}\]

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{-0.28662x^{-1.1}+1.7x^{-2}-0.843x^{1.1}+0.5x^{0.2}-(18.546x^{1.1}-1.1x^{0.2})}{(3.4+x^{2.2})^2}\]

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{-0.28662x^{-1.1}+1.7x^{-2}-(0.843+18.546)x^{1.1}+(0.5+1.1)x^{0.5}}{(3.4+x^{2.2})^2}\]

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{-0.28662x^{-1.1}+1.7x^{-2}-19.389x^{1.1}+1.6x^{0.5}}{(3.4+x^{2.2})^2}\]

myininaya (myininaya):

if you really want to "simplify more" you need to multiply top and bottom by \[x^{3.1}\]

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{-0.28662x^{2}+1.7x^{1.1}-19.389x^{4.2}+1.6x^{3.6}}{x^{3.1}(3.4+x^{2.2})^2}\]

OpenStudy (jamesj):

Gruesome. Kudos to you myininaya.

myininaya (myininaya):

i didn;t want to do it lol

OpenStudy (jamesj):

I really don't like these sorts of calculations, so messy!

myininaya (myininaya):

there is no way i'm checking my work i will leave that to you riley

myininaya (myininaya):

lol sorry

OpenStudy (riley):

I'm trying to work back through it too, taking me a few minutes though . . . just a sec >_< And thank you for working through it like that myininaya, I know it's very time consuming, so thank you.

myininaya (myininaya):

well i don't like doing these types because a mistake is so easy to make

myininaya (myininaya):

if i was your instructor i would not have asked you to simplify something like this

myininaya (myininaya):

likes james said this is a monster

OpenStudy (riley):

I know, I don't think anyone in my class has gotten this problem correct yet. I have to switch rooms though, but I'm gonna go to the computer lab and work on homework. So I will be right back. Almost done working through it, I keep messing up though . . . uuughh. Be right back. >_<

OpenStudy (riley):

And thank you again Myininaya, very very very much.

myininaya (myininaya):

np lol

OpenStudy (riley):

Lol, you got it correct Myininaya!! It took me a little while to work through it all, and I had to go back and check over everything a bunch, but in the end I came up with the same thing you did (finally) and I put that answer in and it's correct!!! Thank you so very very very very much Myininaya; I think thats the toughest problem I've had to do all year, and your help was VERY much appreciated. *hugs* <3